A set of 50 data values has a mean of 15 and a variance of 36.
Find the standard score of a data value = 30.


Answer by stanbon(75887)

A set of 50 data values has a mean of 15 and a variance of 36.
Find the standard score of a data value = 30.
-----
z(30) = (30-15)/6 = 15/6
===============================
Cheers,
stan H.